# FollowIR: Evaluating and Teaching Information Retrieval Models to Follow Instructions

### v1

@@ 3.1 Evaluation Metrics for FollowIR
Our benchmark provides two ways of measuring instruction following: (1) standard retrieval metrics when using the instructions with the queries and (2) pairwise evaluation of instruction following. (1) We use typical IR evaluation metrics but use the instruction along with the query: these metrics are mean average precision (MAP) for Core17/Robust04 and normalized discounted cumulative gain at 5 (nDCG@5) for News21. For (2) we use our novel pairwise evaluation metric that measures the delta in scores when following the modified instructions instead of the original. [^5] Note that we do not show standard retrieval results on the modified instruction’s relevant document set, as standard retrieval scores cannot be directly compared across different query relevance annotations (qrels).
Our new pairwise evaluation metric, $p$-MRR, measures rank-wise changes between queries. In developing this metric we had the following desiderata: it should compare the results of the original instruction to those of the new instruction, it should have a standardized range from worst possible change in instruction-following score (i.e. -1) to best possible instruction-following score (i.e. 1) with an option for no change when using different instructions (i.e. 0), and finally should take into account the document rank so that changes from rank 1 to rank 2 are more prominent than changes from rank 99 to 100. Given the above qualifications, we use the following equation applied to each relevant document per query (where MRR is mean reciprocal rank, $R_{og}$ is the rank of the document when using the original instruction and $R_{new}$ is the new rank):
@@
For the final score, we average first within a given query and then over all queries in the corpora – i.e. macro-averaging across queries, to handle the different number of relevant documents per query.
@@ 4 Evaluating Instruction Following
In this section we describe the models we evaluate, results of those models on FollowIR, and ablations performed to better understand the behavior of current models.
@@ 4.1 Evaluation Settings
Instructions in IR Training
Most retrieval models using instructions received roughly one instruction per retrieval dataset, which generally defined the domain (e.g. “Financial"), document size (sentence, passage, etc.), and task format. This includes INSTRUCTOR models (Su et al., 2022), the bi-encoder TART model trained from Contriever (Asai et al., 2022), the reranker TART trained from FLAN-T5 (Chung et al., 2022), E5 Mistral-Instruct (Wang et al., 2023), and GritLM (Muennighoff et al., 2024). We also include BGE models (Xiao et al., 2023) in this category, although they are trained with only one instruction total for each broad task (retrieval, clustering, etc.).
API Models
We use two of the best performing API embedding models: Cohere’s v3 English and OpenAI’s Text-Embedding-v3-Large. It is unknown what these model’s training procedures were – including if they were trained on instructions or not – thus we place them in a distinct category.
Instruction-Tuned LLMs
We also evaluate several instruction-tuned LLMs to be used as rerankers, including FLAN-T5 (Chung et al., 2022), Llama v2 (Touvron et al., 2023b), and Mistral-Instruct-v0.2 (Jiang et al., 2023). We evaluate these models in the same fashion as MonoT5 rerankers, comparing the true and false tokens. Note that these models were not trained on any retrieval-specific data.
@@
No-Instruction IR
BM25 12.2 -2.8 21.3 +2.5 8.1 -0.7 13.9 -0.3
E5-base-v2 14.5 -6.8 21.6 -4.1 14.0 -2.9 16.7 -4.6
E5-large-v2 18.1 -4.1 24.9 -2.2 17.0 +0.1 20.0 -2.1
Contriever 20.3 -6.1 24.0 -1.8 15.3 -2.5 19.9 -3.5
MonoBERT 21.5 -9.7 26.3 -4.4 18.4 -1.3 22.1 -5.1
MonoT5-base 16.3 -5.8 11.9 -1.2 12.2 -3.5 13.5 -3.5
MonoT5-3B 27.8 +5.6 18.6 +7.5 18.1 +1.7 21.5 +4.9
Instruction-IR
BGE-base 17.5 -6.4 23.8 -0.2 14.6 -2.7 18.6 -3.1
BGE-large 18.1 -7.8 26.4 +0.1 15.0 +0.1 19.8 -2.5
TART-Contriever 14.1 -7.8 21.9 +0.0 12.4 -1.3 16.1 -3.0
INSTRUCTOR-base 14.4 -5.6 16.3 -2.5 14.7 -2.2 15.1 -3.4
INSTRUCTOR-xl 15.5 -2.1 14.6 -4.3 14.4 -0.6 14.8 -2.3
TART-FLAN-T5-xl 25.2 -0.8 20.3 -1.1 17.0 +2.8 20.8 +0.3
GritLM-7B 29.0 -1.4 25.2 +2.1 20.8 +2.6 25.0 +1.1
APIs
Cohere v3 English 22.9 -3.3 23.6 -3.1 20.6 +2.7 22.4 -1.2
OpenAI v3 Large 27.9 -5.7 30.0 -3.3 21.4 -0.2 26.4 -3.1
Instruct LMs
FLAN-T5-base 6.8 +5.0 2.2 +1.1 6.5 -3.2 5.2 +1.0
FLAN-T5-large 15.1 +4.0 8.5 +7.7 11.5 +1.2 11.7 +4.3
Llama-2-7B-chat 6.9 +1.6 13.3 +2.1 5.4 +3.6 8.5 +2.4
Mistral-7B-instruct 24.1 +12.2 22.9 +10.5 19.6 +13.4 22.2 +12.0
FollowIR-7B 25.9 +13.6 25.7 +10.8 20.0 +16.3 23.9 +13.6
Table 3: Evaluating instruction-following on FollowIR. $p$-MRR is a new pairwise evaluation metric measuring instruction following when instructions change, ranging from -100 to 100 (higher is better). We see that the only models that show any success at following instructions are large models (3B+ parameters) or instruction-tuned LLMs that haven’t been trained on retrieval tasks.
@@ 4.2 Results
We see that the no-instruction models range in standard IR metrics (in terms of nDCG@5 and MAP) but generally have negative scores for $p$-MRR (up to -9.7). The only non-instruction model to score positively on average is MonoT5-3B.
@@
We again see that these models have generally negative scores, with the exception being TART-FLAN-T5-xl (with scores averaging roughly zero) and GritLM which has slightly positive scores for two of the three datasets.
Figure 2: Score difference when changing from no instructions to using instructions formatted as keywords, short text, or the full text. We see that models that can correctly use instructions see gains with the additional information while most other models see decreasing performance as instruction length increases. This indicates that most models struggle with instruction length, using these instructions for keyword search rather than as instructions that define document relevance.
@@
We see that the API models perform strongly in terms of standard IR metrics, with OpenAI’s model performing the highest overall. However, both OpenAI and Cohere’s models perform poorly at instruction-following with negative scores (-3.1 and -2.8 on average, respectively).
Instruct-Tuned LLMs
In contrast to the previous results, all instruction-tuned LLMs show positive results for instruction following, although they have the widest range of performance using standard IR metrics (ranging from very poor scores to some of the higher scores). We see that the best performing model in this category is FollowIR–7B, which we describe in more detail in Section 5.
@@
We see that the only models that show positive results at following instructions are either larger IR models (3B+ parameters) or those that have been explicitly trained to follow instructions (FLAN-T5) without retrieval-specific training data. This aligns with work in the natural language processing community which has shown that the instruction-following ability improves with scale (Brown et al., 2020) and instruct-tuning (Longpre et al., 2023).
@@ 4.3 Analysis
Why do so many models fail to correctly follow instructions when they do well on typical IR metrics such as nDCG and MAP? We answer this question by ablating several components that may impact results: (1) whether IR models are not used to instructions that cannot be used for keyword search and (2) whether they are unused to the length of the longer instruction (as instruction-IR models have been trained on much shorter instructions).
We compare the original query-only result to those where we additionally give the model either the full instruction, a shorter instruction, or keywords from the instruction. We gather these short instructions and keywords by prompting GPT-4-Turbo-1106 to generate them from the original full instruction (for TREC data) or the original short instruction used by the model (for BEIR data). For the full prompt text, please see Appendix E.
These results show that models are (1) using the instruction text as keywords and (2) are unused to longer (with less relevant) text that is in the instructions.
We also confirm that these results hold on datasets outside of TREC collections and show results on three BEIR datasets: SciFact, NFCorpus, and FiQA. We show in Table 4 the original score (using the short instructions from their papers) and the change in score when using just keywords from the instruction (again extracted from GPT-4). We show results only for models which performed poorly for instruction-following. We see that the scores for keywords vs the short instruction are generally similar, with most models seeing a change of around $\pm$ 1 point, except for the strongest instruction-following model, E5-Mistral, seeing a larger drop on some datasets.
Overall
@@
Uses Instruction
TART-Contriever 67.6 -0.3 33.4 -5.3 31.8 -0.4
@@ 5 Teaching Instruction Following
Is it possible to improve model performance in following instructions? We seek to show that fine-tuning on a training set of longer instructions can provide a method for doing so.
We start by gathering a training set to teach models. We collect all TREC narratives (i.e. instructions) from tasks not in FollowIR, consisting of 1836 pairs of queries and narratives. However, we note that this does not provide any positive or negative documents for fine-tuning.
In order to obtain documents for training, we prompt GPT-3.5-Turbo-1106 to generate relevant and not-relevant documents, generating roughly two relevant and non-relevant instances per query. However, these synthetic documents are noisy and contains errors w.r.t. the labels – to remedy this, we perform a round of filtering and use the best performing open-source model from Table 3 (Mistral-7B-Instruct-v0.2) to score each of the generated documents according to the instruction. We then filter the documents according to whether Mistral correctly predicts the generated label, and finally balance the relevant and non-relevant samples, choosing only one relevant and non-relevant document per query. Our total is $\sim$1800 training instances on $\sim$1200 unique query/instructions pairs.
@@
When we evaluate this model on FollowIR (Table 3), we find that the scores consistently improve. Compared to the original Mistral-7B-Instruct-v0.2, our model improves on both standard IR metrics (+7.6% relative improvement) and on instruction following (+13.3% relative improvement). Thus, we can see that it is possible to train IR models to be better instruction followers.
@@ 6 Conclusion
Despite the use of LLMs as the backbone of neural retrieval models, most existing IR models do not take instructions that define document relevance. Further, there is no existing resource that measures how well retrieval models can follow instructions. We build a new benchmark that explicitly measures the instruction following ability of retrieval models and find that nearly all retrieval models do not follow instructions, with the exception of very large models (3B+ parameters) or instruction-tuned LLMs that typically are not used for retrieval. However, we show that it is possible to improve their instruction following ability, and build and release a training corpus for teaching retrieval models to follow instructions. Our new model, FollowIR-7B, shows improvement on both standard retrieval metrics as well as in instruction following. We hope that these resources will allow the community to develop more capable instruction-following retrieval models that can quickly adapt to a relevance definition given flexible natural language text.
@@ 7 Limitations
As evaluating instruction following requires evaluating the documents which changed relevance, we cannot use the full collection for retrieval (as each retriever finds different relevant documents by design). Further, due to licensing restrictions, we cannot distribute the full corpora from the TREC tracks – thus we distribute passages due to fair use laws. However, we show full corpus retrieval results for a subset of models in Appendix C and note similar trends.
@@
Our work is built on the TREC document collections and judgements, as well as new annotation efforts. We do not check for potential errors in the TREC annotations, and our newly gathering annotations may have small errors. Despite these caveats, we see that our dataset still provides a useful evaluation setup for measuring instruction following.
Acknowledgments
Orion Weller is supported by a National Science Foundation Graduate Research Fellowship Program grant. We also thank Marc Marone for their feedback and comments that helped improve this work.
@@ Appendix A Hyperparameters for Fine-Tuning Mistral
We fine-tune Mistral-7B-Instruct-v0.2 using the Mistral template for 8 epochs. We use LoRA and tune the “q_proj,v_proj,o_proj,k_proj" parameters, using a rank of 8 and an alpha of 16. We use bfloat16, a batch size of 32, and a learning rate of 3e-5.
@@ Appendix E Prompts Used
I have instructions that are specific to a question, but I need your help abstracting them to a general task format that I can give to someone else. I need you to turn them into an abstract command that just describe the general abstract task instead (e.g., where the data is from, what the type of document looks like). It is crucial that you read and follow these instructions, you will get a large bonus if you are successful (\$200).
@@
Now you can easily see that the abstract command is vague and describes only a short command about how to get the information you need. Follow this exactly – do not reference specifics (like in the above, "international waters" and "medicare" are not included in the abstract command). You should instead keep the abstract command vague and well, abstract about the task only. Use the word "question".
@@
I have instructions that are specific to a style of retrieval, but I want you to instead just focus on the relevant keywords that are in these instructions. Your job is to return a list of these keywords that are relevant in the query. There are probably one or two relevant keywords to extract only. ## Examples ### Example 1: Instruction: Help me to find a highly related PubMed paper to answer this question. Keywords: ["PubMed"]
### Example 2: Instruction: I want to find an answer for this Trivia question. Can you find some paragraphs that provide evidence from Wikipedia? Keywords: ["Trivia", "Wikipedia"]
### Example 3: Instruction: Check if a Quora question is duplicated with this question. Keywords: ["Quora", "duplicated"]
### Example 4: Instruction: I want to find a related question asked in StackExchange. Can you find one for me? Keywords: ["related", "StackExchange"]
## Your turn Instruction: FILL_TEXT_HERE Keywords (either one or two keywords, that are not "documents", "questions", "answer", or "articles"):
I need you to annotate some data for my business and it is super important that you follow instructions precisely or you will be fired. Given a Google search query and instructions regarding what makes a document relevant, I need you to write two documents: one that would be relevant and one that would not. Search: TITLE_HERE Instructions: NARRATIVE_HERE I need some different options to choose from, so give me three **different** options for both a relevant document and an irrelevant document. They should be **long** paragraph-sized documents ($\sim$300 words each), one on each line. If there is no negation in the instructions, your irrelevant document should be slightly off topic:

### v2

@@ 3.1 Evaluation Metrics for FollowIR
Our benchmark provides two ways of measuring instruction following: (1) standard retrieval metrics when using the instructions with the queries and (2) pairwise evaluation of instruction following. For (1), we use typical IR evaluation metrics but use the instruction along with the query: these metrics are mean average precision (MAP) for Core17/Robust04 and normalized discounted cumulative gain at 5 (nDCG@5) for News21. For (2) we use our novel pairwise evaluation metric that measures the delta in scores when following the modified instructions instead of the original. [^6] Note that we do not show standard retrieval results on the modified instruction’s relevant document set, as standard retrieval scores cannot be directly compared across different query relevance annotations (qrels).
Our new pairwise evaluation metric, $p$-MRR, measures rank-wise changes between queries. In developing this metric we had the following desiderata: it should compare the results of the original instruction to those of the new instruction, it should have a standardized range from worst possible change in instruction-following score (i.e., $-1$) to best possible instruction-following score (i.e., $1$) with an option for no change when using different instructions (i.e., 00), and finally should take into account the document rank so that changes from rank 1 to rank 2 are more prominent than changes from rank 99 to 100. Given the above qualifications, we use the following equation applied to each changed relevance document per query (where MRR is mean reciprocal rank, $R_{og}$ is the rank of the document when using the original instruction and $R_{new}$ is the new rank):
@@
For the final score, we average first within a given query and then over all queries in the corpora—i.e., macro-averaging across queries, to handle the different number of relevant documents per query.
@@ 4 Evaluating Instruction Following
In this section we describe the models we evaluate, their results on FollowIR, and ablations performed to better understand the behavior of current models.
@@ 4.1 Evaluation Settings
No-Instruction IR E5-base-v2 13.4 -6.7 20.9 -2.0 14.0 -2.9 16.1 -3.9
Contriever 19.7 -6.1 22.9 -2.8 15.3 -2.5 19.3 -3.8
MonoBERT 21.0 -9.4 25.1 -0.8 18.4 -0.2 21.5 -3.5
BM25 12.1 -3.1 19.3 -2.1 8.1 -1.1 13.2 -2.1
MonoT5-base 15.7 -6.2 11.0 +5.0 12.2 -4.1 13.0 -1.8
E5-large-v2 17.4 -4.2 24.3 +0.9 17.0 +0.1 19.6 -1.1
MonoT5-3B 27.3 +4.0 16.5 +1.8 18.2 +1.8 20.7 +2.5
Instruction-IR TART-Contriever 14.3 -9.0 21.8 -3.0 13.3 -3.0 16.5 -5.0
INSTRUCTOR-base 17.2 -10.4 22.1 -1.8 15.5 -1.1 18.3 -4.4
E5-mistral 23.1 -9.6 27.8 -0.9 18.3 +0.1 23.1 -3.5
BGE-base 16.8 -6.5 20.0 -0.1 14.6 -2.7 17.1 -3.1
INSTRUCTOR-xl 19.7 -8.1 26.1 -0.9 16.8 +0.7 20.9 -2.8
BGE-large 17.5 -7.8 22.3 +0.6 15.0 +0.1 18.3 -2.4
GritLM-7B 28.6 -1.7 24.4 -1.0 20.8 +2.6 24.6 -0.0
TART-FLAN-T5-xl 24.6 -0.7 12.8 +2.0 17.0 +2.8 18.1 +1.4
APIs OpenAI v3 Large 27.2 -5.8 27.2 -2.0 21.6 -0.2 25.3 -2.7
Cohere v3 English 22.3 -3.6 28.3 +0.2 20.6 +2.8 23.7 -0.2
Google Gecko 23.3 -2.4 29.5 +3.9 23.2 +5.4 25.3 +2.3
Instruct LMs FLAN-T5-base 6.4 +5.3 6.1 -0.1 6.5 -3.3 6.3 +0.6
Llama-2-7B-chat 6.3 +2.0 1.7 +0.2 5.4 +2.8 4.5 +1.7
FLAN-T5-large 14.7 +3.9 8.0 +8.9 11.4 +1.3 11.4 +4.7
GritLM-Reranker 9.7 +6.1 10.2 +3.4 9.8 +8.6 9.9 +6.0
Mistral-7B-instruct 23.2 +12.6 27.2 +4.8 19.7 +13.0 23.4 +10.1
FollowIR-7B 24.8 +13.7 29.6 +6.3 20.0 +16.5 24.8 +12.2
Table 2: Evaluating instruction-following on FollowIR. Introduced in this work, $p$-MRR is a pairwise evaluation metric measuring instruction following when instructions change, ranging from $-100$ to $100$ (higher is better). Generally only models with over 3B parameters or instruction-tuned LMs that haven’t been trained on retrieval tasks show success at following retrieval instruction.
Instructions in IR Training
Most retrieval models using instructions received roughly one instruction per retrieval dataset, which generally defined the domain (e.g., “Financial"), document size (sentence, passage, etc.), and task format. This includes INSTRUCTOR models (Su et al., 2022), the bi-encoder TART model trained from Contriever (Asai et al., 2022), the reranker TART trained from FLAN-T5 (Chung et al., 2022), E5 Mistral-Instruct (Wang et al., 2023a), and GritLM (Muennighoff et al., 2024). We also include BGE models (Xiao et al., 2023) in this category, although they are trained with only one instruction total for each broad task (retrieval, clustering, etc.).
API Models
We use three of the best performing API embedding models: Cohere’s v3 English, Google’s Gecko (Lee et al., 2024) and OpenAI’s Text-Embedding-v3-Large. It is mostly unknown what these models’ training procedures were—including if they were trained on instructions or not—thus we place them in a distinct category. However, we note that Google’s model did explicitly train with instructions, as mentioned in their technical report.
Instruction-Tuned LMs
We also evaluate several instruction-tuned LMs to be used as rerankers, including FLAN-T5 (Chung et al., 2022), Llama v2 (Touvron et al., 2023b), and Mistral-Instruct-v0.2 (Jiang et al., 2023). We evaluate these models in the same fashion as MonoT5 rerankers, comparing the true and false tokens. Note that these models were not trained on any retrieval-specific data.
@@ 4.2 Results
We see that the no-instruction models range widely in standard IR metrics (in terms of nDCG@5 and MAP) but generally have negative scores for $p$-MRR (up to $-3.9$). The only non-instruction model to score positively on average is MonoT5-3B (+2.5 $p$-MRR).
@@
We again see that these models have generally negative scores, with the exception being GritLM (with scores averaging roughly zero) and TART-FLAN-T5-xl which has slightly positive scores for two of the three datasets (with an average of +1.4 $p$-MRR).
Figure 3: Score difference between using no instructions to using instructions formatted as keywords, short text, or the full text. While models that can correctly use instructions see gains with the additional information, most other models see decreasing performance as instruction length increases.
@@
We see that the API models perform strongly in terms of standard IR metrics, with OpenAI’s and Google’s models performing the highest overall. However, Cohere’s and OpenAI’s models perform poorly at instruction-following with negative scores ($-0.2$ and $-2.7$ on average, respectively) whereas Google Gecko has positive scores (+2.3) likely a result of training on more instruction-focused data.
Instruct-Tuned LMs
In contrast to the previous results, all instruction-tuned LMs show positive results for instruction following, although they have the widest range of performance using standard IR metrics (ranging from very poor scores to some of the higher scores). We see that the best performing model in this category is FollowIR–7B, which we describe in more detail in Section 5.
@@
We see that the only models that show positive results at following instructions are either IR models with over 3B parameters or those that have been explicitly trained to follow instructions (e.g. FLAN-T5), without any retrieval-specific supervision. This aligns with work in the natural language processing community which has shown that the instruction-following ability improves with scale (Brown et al., 2020) and supervised instruction-tuning (Longpre et al., 2023).
@@ 4.3 Analysis
Why do so many models fail to correctly follow instructions when they do well on typical IR metrics such as nDCG and MAP? We answer this question by ablating several components that may impact results: (1) whether IR models are not used to text that cannot be used for simple keyword search (i.e. instructions) and (2) whether they are unused to the length of the longer instructions (as current instruction IR models have been trained on much shorter instructions).
@@
Uses Instruction TART-Contriever 67.6 -0.3 33.4 -5.3 31.8 -0.4
@@
To test these, we compare the original query-only result to those where we additionally give the model either the full instruction, a shorter instruction, or keywords from the instruction. We gather these short instructions and keywords by prompting GPT-4-Turbo-1106 to generate them from the original full instruction (for TREC data) or otherwise use the original short instructions given by the authors of the model (for BEIR data). For the full prompt text, please see Appendix E.
These results show that models are (1) using the instruction text as keywords (as performance is higher when using only keywords) and (2) are not able to utilize the extra information in the instructions (as they generally decrease in performance with this additional information).
We also confirm that these results hold on datasets outside of TREC collections and show results on three BEIR datasets: SciFact, NFCorpus, and FiQA. We show in Table 3 the original score (using the short instructions from their papers) and the change in score when using just keywords from the instruction (again extracted from GPT-4). We show results only for models which performed poorly for instruction-following. We see that the scores for keywords vs the short instruction are generally similar, with most models seeing a change of around $\pm$ 1 point, except for the strongest of the non-instruction-following models, E5-Mistral, seeing a larger drop on some datasets.
Overall
@@ 5 Teaching Instruction Following
Is it possible to improve model performance in following instructions? We show that fine-tuning on a training set of longer instructions can provide a method for doing so.
We start by gathering a training set to teach models. We collect all TREC narratives (i.e., instructions) from tasks not in FollowIR, consisting of 1836 pairs of queries and narratives. However, we note that this does not provide any positive or negative documents for fine-tuning.
In order to obtain documents for training, we prompt GPT-3.5-Turbo-1106 to generate relevant and not-relevant documents, generating roughly two relevant and non-relevant instances per query. However, these synthetic documents are noisy and contains errors w.r.t. the labels—to remedy this, we perform a round of filtering and use the best performing open-source model from Table 2 (Mistral-7B-Instruct-v0.2) to score each of the generated documents according to the instruction. We then filter the documents according to whether Mistral correctly predicts the generated label, and finally balance the relevant and non-relevant samples, choosing only one relevant and non-relevant document per query. Our total is $\sim$1800 training instances on $\sim$1200 unique query/instructions pairs.
@@
Model Robustness@10
BM25 26.9
TART-Contriever 47.5
RepLLaMa 52.6
E5-Mistral 55.4
Mistral-7B-instruct 35.3
FollowIR-7B 71.5
Figure 4: Performance on the InstructIR benchmark using their “Robustness@10" scores, e.g. the min nDCG@10 score across 10 instructions. Upper portion is bi-encoders while lower is rerankers.
When we evaluate this model on FollowIR (Table 2), we find that the scores consistently improve. Compared to the original Mistral-7B-Instruct-v0.2, our model improves on both standard IR metrics (+6.0% relative improvement) and on instruction following (+20.8% relative improvement). We also show that this improvement holds on the concurrent InstructIR dataset (Table 4), where FollowIR-7B scores double the base Mistral-7B scores (71.5 Robustness@10 vs 35.3) and is the top scoring model overall. Thus, we can see that it is possible to train IR models to be better instruction followers.
@@ 6 Conclusion
Despite the use of LMs as the backbone of neural retrieval models, most existing IR models do not take instructions that define document relevance. Further, there is no existing resource that measures how well retrieval models can follow instructions. We build a new benchmark that explicitly measures the instruction following ability of retrieval models and find that nearly all retrieval models do not follow instructions, with the exception of larger models (3B+ parameters) or instruction-tuned LMs that typically are not used for retrieval. However, we show that it is possible to improve their instruction following ability, and build and release a training corpus for teaching retrieval models to follow instructions. Our new model, FollowIR-7B, shows improvement on both standard retrieval metrics as well as in instruction following. We hope that these resources will allow the community to develop more capable instruction-following retrieval models that can quickly adapt to a relevance definition given flexible natural language text.
@@ 7 Limitations
As our setup for evaluating instruction following requires evaluating the documents which changed relevance, we cannot use the full collection for retrieval (as each retriever finds different relevant documents by design). Further, due to licensing restrictions, we cannot distribute the full corpora from the TREC tracks—thus we distribute passages due to fair use laws. However, we show full corpus retrieval results for a subset of models in Appendix C and note similar trends in terms of the lack of instruction following.
@@
Our work is built on the TREC document collections and judgements, as well as new annotation efforts. We do not check for potential errors in the TREC annotations, and our newly gathered annotations may have small errors. Despite these caveats, we see that our dataset still provides a useful evaluation setup for measuring instruction following.
@@ Appendix E Prompts Used
Synthetic Document Creation I need you to annotate some data for my business and it is super important that you follow instructions precisely or you will be fired. Given a Google search query and instructions regarding what makes a document relevant, I need you to write two documents: one that would be relevant and one that would not. Search: TITLE_HERE Instructions: NARRATIVE_HERE I need some different options to choose from, so give me three **different** options for both a relevant document and an irrelevant document. They should be **long** paragraph-sized documents ($\sim$300 words each), one on each line. If there is no negation in the instructions, your irrelevant document should be slightly off topic:
Short Instructions to Keywords I have instructions that are specific to a style of retrieval, but I want you to instead just focus on the relevant keywords that are in these instructions. Your job is to return a list of these keywords that are relevant in the query. There are probably one or two relevant keywords to extract only. ## Examples
### Example 1:
Instruction: Help me to find a highly related PubMed paper to answer this question.
Keywords: ["PubMed"]
### Example 2:
Instruction: I want to find an answer for this Trivia question. Can you find some paragraphs that provide evidence from Wikipedia?
Keywords: ["Trivia", "Wikipedia"]
### Example 3:
Instruction: Check if a Quora question is duplicated with this question.
Keywords: ["Quora", "duplicated"]
### Example 4:
Instruction: I want to find a related question asked in StackExchange. Can you find one for me?
Keywords: ["related", "StackExchange"]
## Your turn
Instruction: FILL_TEXT_HERE
Keywords (either one or two keywords, that are not "documents", "questions", "answer", or "articles"):
Full Instructions to Short Instructions I have instructions that are specific to a question, but I need your help abstracting them to a general task format that I can give to someone else. I need you to turn them into an abstract command that just describe the general abstract task instead (e.g., where the data is from, what the type of document looks like). It is crucial that you read and follow these instructions, you will get a large bonus if you are successful (\$200).
@@
Now you can easily see that the abstract command is vague and describes only a short command about how to get the information you need. Follow this exactly—do not reference specifics (like in the above, "international waters" and "medicare" are not included in the abstract command). You should instead keep the abstract command vague and well, abstract about the task only. Use the word "question".